# Computing Inconsistency Measures Under Differential Privacy

### v2

@@ 2.1. Database and Constraints
[…]
@@
[…]
@@ 2.3. Differential Privacy
Minimal inconsistency measure $\mathcal{I_{\mathsf{MI}}}$ the no. of edges $O(|\Sigma|n^{2})$ $n$ $O(|\Sigma|n^{2}+|\Theta|m)$ $-\tilde{q}_{{\text{opt}}}(D,\varepsilon_{2})+O(\frac{\theta_{\max}\ln|\Theta|}{\varepsilon_{1}})$
Problematic measure $\mathcal{I_{\mathsf{P}}}$ the no. of nodes with positive degrees $O(|\Sigma|n^{2})$ $n$ $O(|\Sigma|n^{2}+|\Theta|m)$ $-\tilde{q}_{{\text{opt}}}(D,\varepsilon_{2})+O(\frac{\theta_{\max}\ln|\Theta|}{\varepsilon_{1}})$
@@ 3. Inconsistency Measures under DP
Inadequacy of $\mathcal{I_{\mathsf{D}}}$ and $\mathcal{I_{\mathsf{MC}}}$. We note that two inconsistency measures are less suitable for DP. First, the drastic measure $\mathcal{I_{\mathsf{D}}}$ is a binary measure that outputs $1$ if at least one conflict exists in the dataset and 00 otherwise. Due to its binary nature, the measure’s sensitivity is 1, meaning adding or removing a single row can significantly alter the result. Adding DP noise to such a binary measure can render it meaningless.
One way to compute the $\mathcal{I_{\mathsf{D}}}$ measure could be to consider a proxy of $\mathcal{I_{\mathsf{D}}}$ by employing a threshold-based approach that relies on $\mathcal{I_{\mathsf{P}}}$ or $\mathcal{I_{\mathsf{MI}}}$. For example, if these measures are below a certain given number, we return 00 and, otherwise, $1$. A recent work (Patwa et al., 2023) addresses similar problems for synthetic data by employing the exponential mechanism. However, since we focus on directly computing the measures under DP, we leave this intriguing subject for future work.
@@ 4.1. Graph Projection Approach for $\mathcal{I_{\mathsf{MI}}}$ and $\mathcal{I_{\mathsf{P}}}$
A primary utility challenge in achieving node-DP for graph statistics is their high sensitivity. In the worst case, removing a single node from a graph of $n$ nodes can result in removing $(n-1)$ edges. To mitigate this issue, the state-of-the-art approach (Day et al., 2016) first projects the graph $\mathcal{G}$ onto a $\theta$-bounded graph $\mathcal{G}_{\theta}$, where the maximum degree is no more than $\theta$. Subsequently, the edge count of the transformed graph is perturbed by the Laplace mechanism with a sensitivity value of less than $n$. However, the choice of $\theta$ is critical for accurate estimation. A small $\theta$ reduces Laplace noise due to lower sensitivity, but results in significant edge loss during projection. Conversely, a $\theta$ close to $n$ preserves more edges but increases the Laplace noise. Prior work addresses this balance using the exponential mechanism (EM) to prefer a $\theta$ that minimizes the combined errors arising from graph projection and the Laplace noise.
@@ 1 Find the maximum value in $\Theta$ as $\theta_{\max}$
3 Sample $\theta^{*}$ with prob $\propto\exp(\frac{\varepsilon_{1}q_{\varepsilon_{2}}(\mathcal{G},\theta_{i})}{2\theta_{\max}})$
@@
1 2 $\sqrt{2}$ $-2-\sqrt{2}$
2 1 $2\sqrt{2}$ $-1-2\sqrt{2}$
3 0 $3\sqrt{2}$ $-3\sqrt{2}$
@@
For $\mathcal{I_{\mathsf{P}}}$, we can analyze a worst-case scenario where the graph is a star with $n$ nodes such that there is an internal node connected to all other $n-1$ nodes, and the threshold $\theta$ for edge addition is $n$. The edge addition algorithm would play a minimal role, and no edges would be truncated. For a neighboring graph that differs on the internal node, all edges of the graph are removed (connected to the internal node), and the $\mathcal{I_{\mathsf{P}}}=0$ (no problematic nodes), making the sensitivity for $\mathcal{I_{\mathsf{P}}}$ in this worst-case $=n$.
@@
This theorem indicates that the error incurred by Algorithm 2 with Algorithm 3 is directly proportional to the log of the candidate size $|\Theta|$ and the sensitivity of the quality function. The $\beta$ parameter in the theorem is a controllable probability parameter. According to the accuracy requirements of a user’s analysis, one may set $\beta$ as any value less than this upper bound. For example, if we set $\beta=0.01$, then our theoretical analysis of Algorithm 2 that says the algorithm’s output being close to the true answer will hold with a probability of $0.991-\beta=$ 0.99. We also show a plot to show the trend of the utility analysis as a function of $\beta$ in Appendix A.5 (author, s). Without prior knowledge about the graph, $\theta_{\max}$ is usually set as the number of nodes $n$, and $\Theta$ includes all possible degree values up to $n$, resulting in poor utility. Fortunately, for our use case, the edges in the graph arise from the DCs that are available to us. In the next section, we show how we can leverage these constraints to improve the utility of our algorithm by truncating candidates in the set $\Theta$.
@@ 4.2. Optimized Parameter Selection
Our experimental study (Section 6) shows that datasets with general DCs have dense conflict graphs, which favors larger $\theta$s for graph projection. Hence, if we learn a small noisy upper bound $\tilde{d}_{{\text{bound}}}$ based on the FDs with LM, we will first prune all degree candidates smaller than $\tilde{d}_{{\text{bound}}}$, but then include $|V|$, which corresponds to the case when no edges are truncated, and Laplace mechanism is applied with the largest possible sensitivity $|V|$, i.e.,
@@
Though the maximum value in $\Theta^{\prime}$ is $|V|$, the sensitivity of the quality function over the candidate set $\Theta^{\prime}$ remains $\tilde{d}_{{\text{bound}}}$. For the $|V|$ candidate, the quality function only depends on the Laplace error $\frac{\sqrt{2}|V|}{\varepsilon_{2}}$ and has no error from $e_{\text{bias}}$ as no edges will be truncated. Despite being tailored for FDs, we show that, in practice, our approach is cheap and performs well for DCs. In Section 6, we show that this approach works well for the dense Adult (Lichman, 2013) dataset where we compute the $\mathcal{I_{\mathsf{P}}}$ using this strategy in Figure 5. Developing a specific strategy for DCs is an important direction of future work.
In practice, one may skip this upper bound calculation process and skip directly to the two-step exponential mechanism if it is known that the graph is too dense or contains few FDs and more general DCs. We discuss this in detail in the experiments section.
@@ 12       Set $\theta_{\max}\leftarrow\theta^{*}$
The utility analysis in Theorem 5 for Algorithm 2 with the basic EM (Algorithm 3) still applies to the optimized EM (Algorithm 4). The basic EM usually has $\theta_{\max}=|D|=|V|$ and the full budget $\varepsilon_{1}$, while the optimized EM has a much smaller $\theta_{\max}$ and slightly lower privacy budget when the graph is sparse. In practice (Section 6), we see significant utility improvements by the optimized EM for sparse graphs. When the graph is dense, we see the utility degrade slightly due to a smaller budget for each EM. However, the degradation is negligible with respect to the true inconsistency measure.
@@ 6.1. Experimental Setup
[…]
@@ 6.2. Results
Comparing different strategies for choosing $\theta$. In Figure 5, we present the performance of different algorithm variations in computing $\mathcal{I_{\mathsf{P}}}$ and $\mathcal{I_{\mathsf{MI}}}$ for all datasets using RNnoise at $\alpha=0.01$ and $\varepsilon=1$. The y-axis in each figure shows the logarithmic scaled error, while the x-axis displays the actual measure value, with different colors representing the strategies. The graphs are ordered from most sparse (Stock) to least sparse (Adult) to compare methods for choosing the $\theta$ value at $\varepsilon=1$ ($\varepsilon_{1}=0.4$, $\varepsilon_{2}=0.6$). The methods include all variations described in the algorithm variation section. We note all the strategies are private except baseline 2 (orange dash line) that sets $\theta$ as the true maximum degree of the conflict graph.
@@
Varying privacy budget. Figure 6 illustrates how our algorithms perform at $\varepsilon\in[0.1,0.2,0.5,1.0,2.0,3.0,5.0]$ with varying privacy budgets. The rightmost figure for the repair measure $\mathcal{I_{\mathsf{R}}}$ has a log scale on the y-axis for better readability. We experiment with three datasets of different density properties (sparsest Stock and densest Flight) and show that our algorithm gracefully scales with the $\varepsilon$ privacy budget for all three inconsistency measures. We also observe that the algorithm has a more significant error variation at smaller epsilons ($<1$) except for Stock, which has a larger variation across all epsilons. This happens when the true value for this measure is small, and adding noise at a smaller budget ruins the estimate drastically. For the $\mathcal{I_{\mathsf{R}}}$ measure, the private value reaches with $0.05$ error at $\varepsilon=3$ for Stock and as early as $\varepsilon=0.1$ for others.
@@ A.2.3. Proof for  Lemma 3
Assume, in the worst case, the graph $\mathcal{G}$ is a star graph with $n$ nodes such that there exists an internal node that is connected to all other $n-1$ nodes. In this scenario, there are no nodes that have 0 degrees, and the $\mathcal{I_{\mathsf{P}}}$ measure $=n-0=0$. If the neighbouring graph $\mathcal{G}^{\prime}$ differs on the internal node, all edges of the graph are removed are the $\mathcal{I_{\mathsf{P}}}=n$. The edge addition algorithm $\pi_{\theta}$ would play a minimal role here as $\theta$ could be equal to $n$. ∎

### v3

@@ 2.1. Database and Constraints
We consider a single-relation schema $\mathcal{A}=(A_{1},\ldots,A_{m})$, which is a vector of distinct attribute names $A_{i}$, each associated with a domain ${\tt dom}(A_{i})$ of values. A database $D$ over $\mathcal{A}$ is a associated with a set $\mathit{tids}(D)$ of tuple identifiers, and it maps every identifier $i\in\mathit{tids}(D)$ to a tuple $D[i]=(a_{1},\dots,a_{m})$ in $A_{1}\times\dots\times A_{m}$. A database $D^{\prime}$ is a subset of $D$, denoted $D^{\prime}\subseteq D$, if $D^{\prime}$ is obtained from $D$ by deleting zero or more tuples, that is, $\mathit{tids}(D^{\prime})\subseteq\mathit{tids}(D)$ and $D^{\prime}[i]=D[i]$ for all $i\in\mathit{tids}(D^{\prime})$.
Following previous work on related topics (Ge et al., 2021a; Livshits et al., 2020b), we focus on Denial Constraints (DCs) on pairs of tuples. Using the formalism of Tuple Relational Calculus, such a DC is of the form $\forall t,t^{\prime}\neg\big{(}\varphi_{1}\land\dots\land\varphi_{k}\big{)}$ where each $\varphi_{j}$ is a comparison $\sigma_{1}\circ\sigma_{2}$ so that: (a) each of $\sigma_{1}$ and $\sigma_{2}$ is either $t[A_{i}]$, or $t^{\prime}[A_{i}]$, or $a$, where $A_{i}$ is some attribute and $a$ is a constant value, and (b) the operator $\circ$ belongs to set ${<,>,,,=,}\{$$\leq$$\geq,$$\neq\}$ of comparisons. This DC states that there cannot be two tuples $t$ and $t^{\prime}$ such that all comparisons $\varphi_{j}$ hold true (i.e., at least one $\varphi_{j}$ should be violated).
@@
A common way of capturing the violations of $\Sigma$} in $D$ is through the conflict graph $\mathcal{G}_{\Sigma}^{D}$}, which is the graph $(V,E)$, where $V=\mathit{tids}(D)$ an edge $e=\mathord{\{i,j\}}\in E$ occurs whenever the tuples $D[i]$ and jointly $D[j]$ violate $\Sigma$}. To simplify the notation, we may write simply $\mathcal{G}$ instead of $\mathcal{G}_{\Sigma}^{D}$} when there is no risk of ambiguity.
@@ 2.3. Differential Privacy
Minimal inconsistency measure $\mathcal{I_{\mathsf{MI}}}$ the no. of edges $O(|\Sigma|n^{2})$ $n$ $O(|\Sigma|n^{2}+|\Theta|m)$ $\tilde{q}_{{\text{opt}}}(D,\varepsilon_{2})+O(\frac{\theta_{\max}\ln|\Theta|}{\varepsilon_{1}})$
Problematic measure $\mathcal{I_{\mathsf{P}}}$ the no. of nodes with positive degrees $O(|\Sigma|n^{2})$ $n$ $O(|\Sigma|n^{2}+|\Theta|m)$ $\tilde{q}_{{\text{opt}}}(D,\varepsilon_{2})+O(\frac{\theta_{\max}\ln|\Theta|}{\varepsilon_{1}})$
@@ 3. Inconsistency Measures under DP
Inadequacy of $\mathcal{I_{\mathsf{D}}}$ and $\mathcal{I_{\mathsf{MC}}}$. We note that two inconsistency measures are less suitable for DP. First, the drastic measure $\mathcal{I_{\mathsf{D}}}$ is a binary measure that outputs $1$ if at least one conflict exists in the dataset and 0 otherwise. Due to its binary nature, the measure’s sensitivity is 1, meaning adding or removing a single row can significantly alter the result. Adding DP noise to such a binary measure can render it meaningless.
One way to compute the $\mathcal{I_{\mathsf{D}}}$ measure could be to consider a proxy of $\mathcal{I_{\mathsf{D}}}$ by employing a threshold-based approach that relies on $\mathcal{I_{\mathsf{P}}}$ or $\mathcal{I_{\mathsf{MI}}}$. For example, if these measures are below a certain given number, we return 0 and, otherwise, $1$. A recent work (Patwa et al., 2023) addresses similar problems for synthetic data by employing the exponential mechanism. However, since we focus on directly computing the measures under DP, we leave this intriguing subject for future work.
@@ 4.1. Graph Projection Approach for $\mathcal{I_{\mathsf{MI}}}$ and $\mathcal{I_{\mathsf{P}}}$
A primary utility challenge in achieving node-DP for graph statistics is their high sensitivity. In the worst case, removing a single node from a graph of $n$ nodes can result in removing (n1)(n-1) edges. To mitigate this issue, the state-of-the-art approach (Day et al., 2016) first projects the graph $\mathcal{G}$ onto a $\theta$-bounded graph $\mathcal{G}_{\theta}$, where the maximum degree is no more than $\theta$. Subsequently, the edge count of the transformed graph is perturbed by the Laplace mechanism with a sensitivity value of less than $n$. However, the choice of $\theta$ is critical for accurate estimation. A small $\theta$ reduces Laplace noise due to lower sensitivity, but results in significant edge loss during projection. Conversely, a $\theta$ close to $n$ preserves more edges but increases the Laplace noise. Prior work addresses this balance using the exponential mechanism (EM) to prefer a $\theta$ that minimizes the combined errors arising from graph projection and the Laplace noise.
@@ 1 Find the maximum value in $\Theta$ as $\theta_{\max}$
3
4Sample $\theta^{*}$ with prob $\propto\exp(\frac{\varepsilon_{1}q_{\varepsilon_{2}}(\mathcal{G},\theta_{i})}{2\theta_{\max}})$
@@
1 2 $\sqrt{2}$ $2-\sqrt{2}$
2 1 $2\sqrt{2}$ $1-2\sqrt{2}$
3 0 $3\sqrt{2}$ $3\sqrt{2}$
@@
For $\mathcal{I_{\mathsf{P}}}$, we can analyze a worst-case scenario where the graph is a star with $n$ nodes such that there is an internal node connected to all other n1n-1 nodes, and the threshold $\theta$ for edge addition is $n$. The edge addition algorithm would play a minimal role, and no edges would be truncated. For a neighboring graph that differs on the internal node, all edges of the graph are removed (connected to the internal node), and the $\mathcal{I_{\mathsf{P}}}=0$ (no problematic nodes), making the sensitivity for $\mathcal{I_{\mathsf{P}}}$ in this worst-case =n=n.
@@
This theorem indicates that the error incurred by Algorithm 2 with Algorithm 3 is directly proportional to the log of the candidate size $|\Theta|$ and the sensitivity of the quality function. The $\beta$ parameter in the theorem is a controllable probability parameter. According to the accuracy requirements of a user’s analysis, one may set $\beta$ as any value less than this upper bound. For example, if we set $\beta=0.01$, then our theoretical analysis of Algorithm 2 that says the algorithm’s output being close to the true answer will hold with a probability of $1-\beta=0.99$. We also show a plot to show the trend of the utility analysis as a function of $\beta$ in Appendix A.5 (author, s). Without prior knowledge about the graph, $\theta_{\max}$ is usually set as the number of nodes $n$, and $\Theta$ includes all possible degree values up to $n$, resulting in poor utility. Fortunately, for our use case, the edges in the graph arise from the DCs that are available to us. In the next section, we show how we can leverage these constraints to improve the utility of our algorithm by truncating candidates in the set $\Theta$.
@@ 4.2. Optimized Parameter Selection
[…]
@@
Though the maximum value in $\Theta^{\prime}$ is $|V|$, the sensitivity of the quality function over the candidate set $\Theta^{\prime}$ remains $\tilde{d}_{{\text{bound}}}$. For the $|V|$ candidate, the quality function only depends on the Laplace error $\frac{\sqrt{2}|V|}{\varepsilon_{2}}$ and has no error from $e_{\text{bias}}$ as no edges will be truncated. Despite being tailored for FDs, we show that, in practice, our approach is cheap and performs well for DCs. In Section 6, we show that this approach works well for the dense Adult (Lichman, 2013) dataset where we compute the $\mathcal{I_{\mathsf{P}}}$ using this strategy in Figure 5. Developing a specific strategy for DCs is an important direction of future work. In practice, one may skip this upper bound calculation process and skip directly to the two-step exponential mechanism if it is known that the graph is too dense or contains few FDs and more general DCs. We discuss this in detail in the experiments section.
@@ 12   Set $\theta_{\max}\leftarrow\theta^{*}$
The utility analysis in Theorem 5 for Algorithm 2 with the basic EM (Algorithm 3) still applies to the optimized EM (Algorithm 4). The basic EM usually has $\theta_{\max}=|D|=|V|$ and the full budget $\varepsilon_{1}$, while the optimized EM has a much smaller $\theta_{\max}$ and slightly lower privacy budget when the graph is sparse.
In practice (Section 6), we see significant utility improvements by the optimized EM for sparse graphs. When the graph is dense, we see the utility degrade slightly due to a smaller budget for each EM. However, the degradation is negligible with respect to the true inconsistency measure.
@@ 6.1. Experimental Setup
    • If $\circ\in\{=,\leq,\geq\}$, change either $t_{i}[a_{1}]$ or $t_{j}[a_{2}]$ or vice versa (the choice is random).
    • If $\circ\in\{<,>,\neq\}$, change either $t_{i}[a_{1}]$ or $t_{2}[a_{2}]$ (the choice is again random) to another value from the active domain of the attribute such that $\phi$ is satisfied, if such a value exists, or a random value in the appropriate range otherwise.
@@ 6.2. Results
Comparing different strategies for choosing $\theta$. In Figure 5, we present the performance of different algorithm variations in computing $\mathcal{I_{\mathsf{P}}}$ and $\mathcal{I_{\mathsf{MI}}}$ for all datasets using RNnoise at $\alpha=0.01$ and $\varepsilon=1$. The y-axis in each figure shows the logarithmic scaled error, while the x-axis displays the actual measure value, with different colors representing the strategies. The graphs are ordered from most sparse (Stock) to least sparse (Adult) to compare methods for choosing the $\theta$ value at $\varepsilon=1$ ($\varepsilon_{1}=0.4$, $\varepsilon_{2}=0.6$). The methods include all variations described in the algorithm variation section.
We note all the strategies are private except baseline 2 (orange dash line) that sets $\theta$ as the true maximum degree of the conflict graph.
@@
Varying privacy budget. Figure 6 illustrates how our algorithms perform at $\varepsilon\in[0.1,0.2,0.5,1.0,2.0,3.0,5.0]$ with varying privacy budgets. The rightmost figure for the repair measure $\mathcal{I_{\mathsf{R}}}$ has a log scale on the y-axis for better readability. We experiment with three datasets of different density properties (sparsest Stock and densest Flight) and show that our algorithm gracefully scales with the $\varepsilon$ privacy budget for all three inconsistency measures. We also observe that the algorithm has a more significant error variation at smaller epsilons (<1<1) except for Stock, which has a larger variation across all epsilons. This happens when the true value for this measure is small, and adding noise at a smaller budget ruins the estimate drastically. For the $\mathcal{I_{\mathsf{R}}}$ measure, the private value reaches with $0.05$ error at $\varepsilon=3$ for Stock and as early as $\varepsilon=0.1$ for others.
@@ A.2.3. Proof for  Lemma 3
Assume, in the worst case, the graph $\mathcal{G}$ is a star graph with $n$ nodes such that there exists an internal node that is connected to all other n1n-1 nodes. In this scenario, there are no nodes that have 0 degrees, and the $\mathcal{I_{\mathsf{P}}}$ measure =n0=0=n-0=0. If the neighbouring graph $\mathcal{G}^{\prime}$ differs on the internal node, all edges of the graph are removed are the $\mathcal{I_{\mathsf{P}}}=n$. The edge addition algorithm $\pi_{\theta}$ would play a minimal role here as $\theta$ could be equal to $n$. ∎